OpenStudy (anonymous):

Given the function f(x) = log2(x + 6), find the value of f−1(3).

OpenStudy (anonymous):

f−1(3) = 2 f−1(3) = 3 f−1(3) = 9 f−1(3) = 18

OpenStudy (anonymous):

Your answer is the last choice